OpenStudy (anonymous):

14. What is 512? Use mental math. (1 point) 2,500 2,401 2,601 2,551

OpenStudy (anonymous):

Makes not sense to me.

OpenStudy (anonymous):

doesn't make sense

sam (.sam.):

Maybe we have mental prob, lol (joking)

OpenStudy (lgbasallote):

that's why it says use mental math...we need to predict the question >:))

OpenStudy (anonymous):

This question was asked before at openstudy: http://openstudy.com/study#/updates/4f5432ffe4b01ed61383314e

OpenStudy (anonymous):

51^2

OpenStudy (anonymous):

Lol :D

sam (.sam.):

2601

OpenStudy (anonymous):

2601

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

lol howq did you find that fool that was a month ago

OpenStudy (anonymous):

\[51^2= (50+1)^2 = 2500+100+1=2601 \]